HUBBLE SPACE TELESCOPE DAILY REPORT #2167 PERIOD COVERED: 0000Z (UTC) 07/16/98 - 0000Z (UTC) 07/17/98 Daily Status Report as of 198/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed NIC/2 7853 (Deep IR Imaging of Two Gas-Rich Radio Galaxies) The Near Infrared Camera (Camera Two) was used to make high resolution continuum, line and polarization images of the gas-rich radio galaxy 3C305. This is to obtain a deep IR image to locate the true nuclei and clarify the structure of the galaxies in order to test merger models. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Three Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.4 Completed STIS/CCD 7910 (STIS Extragalactic Parallel Survey) The Space Telescope Imaging Spectrograph (CCD) was used to make extragalactic slitless spectrographic observations as part of the parallel high galactic latitude survey to search for emission line galaxies. Data was taken in parallel, and are not subject to proprietary time. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed NIC/1/2/3 7883 (Infrared Imaging of High Redshift {0.4 <= z <= 1} Tully-Fisher Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the high redshifted Tully- Fisher class galaxy GSS-072-74. With the advent of 10 meter class telescopes combined with the high spatial resolution of the HST, it is now possible to study the internal kinematics and structural properties of galaxies out to a redshift of z ~ 1. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Three Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.8 Completed STIS/CCD 8064 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal - Backup) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a non-scripted parallel POMS test. The observations were completed as scheduled, and no problems were reported. 1.9 Completed NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of WF/PC-2 7616 (Rapid Response to Anomalous Activity in Jupiter's Atmosphere) The WF/PC-2 was used to make observations of anomalous activity in the atmosphere of the planet Jupiter. The well-documented record of aperiodic disturbances within Jupiter's atmosphere reveals that their onset is marked by an intensely white cloud that expands and interacts with the local winds. The observations were completed as planned, and no problems were reported. 1.11 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed NIC/1/2/3 7896 (A NIR Snapshot Survey of Ultraluminous IR Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make snapshot survey observations of the ultraluminous infrared galaxy IR20087-03. Nearly all of these galaxies show evidence for a recent tidal interaction, and we will search for evidence of the merger progenitors near the center of each galaxy, a task made significantly easier with NIR imaging. The observations were completed as planned, and no anomalies were noted. 1.13 Completed NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around the main sequence star TWA8B. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.14 Completed WF/PC-2/NIC/1/2/STIS/CCD/MA2 7560 (STIS GTO Observations) The WF/PC-2, the Near Infrared Camera (Cameras One and Two) and the Space Telescope Imaging Spectrograph (CCD and MA2) were used to make observations of galaxy NGC-3995. The solar blind and long slit capabilities of STIS will be used to investigate the population of hot stars in galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.16 Completed NIC/1/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera One and Two) was used to make a search for low mass/sub-luminous companions to the low mass M class star LHS3003. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.17 Completed NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy NGC-4636 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.18 Completed FOC/96 7923 (FOC F/96 Turnon) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: The recharge of Battery 1 continues. Using ROP SR-1, the SSR EDAC error counter was cleared four times. The NSSC-1 status buffer was dumped and the pointer was reset at 198/0049Z per ROP NS-3. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH